OpenStudy (anonymous):

If E and F are mutually exclusive events, then P (E ∪ F) = P (E) + P (F).in this case , will P (E) + P (F) BE EQUAL TO 1? PLS EXPLAIN WITH REASON.

OpenStudy (anonymous):

@hba

OpenStudy (unklerhaukus):

the sum of all probabilities is one

OpenStudy (anonymous):

thank u @UnkleRhaukus

OpenStudy (unklerhaukus):

lets consider some mutually exclusive events , say i flip a coin , the possible events are heads and tails, and these events are mutually exclusive, the probability of each event is 50% the probability of heads OR tails is P(h+t)=P(h) +P(t) = 50% + 50% = 100% =1